Determine and Use Overhead Rate
Cortez Consulting, Inc., expects the following costs and expenses during the coming year:

Direct labor (@ $27./hr.) $243,000

Sales commissions

55,500
Overhead 216,000

a. Compute a predetermined overhead rate applied on the basis of direct labor hours.

Round to two decimal places, if applicable.

$Answer

b. Prepare a general journal entry to apply overhead during an interim period when 2,250 direct labor hours were worked.

General Journal
Description Debit Credit
AnswerConsulting overheadCost of goods soldFinished goods inventoryWork in process inventory Answer Answer
AnswerConsulting overheadCost of goods soldFinished goods inventoryWork in process inventory Answer Answer

c. What amount of overhead would be assigned to Job 466, to which $270 in direct labor had been charged?

Round to two decimal places, if applicable.

$Answer
Answer #1

a)

Direct labor cost = $243,000

Direct labor rate = $27 per hour

Direct labor hours = Direct labor cost/Direct labor rate

= 243,000/27

= 9,000

Predetermined overhead rate = Estimated overheads/Estimated direct labor hours

= 216,000/9,000

= $24 per direct labor hour

b)

Estimated overheads = Direct labor hours x Predetermined overhead rate

= 2,250 x 24

= $54,000

General Journal
Description Debit Credit
Work in process inventory 54,000
Consulting overhead 54,000

c)

Direct labor hours = Direct labor cost/Direct labor rate

= 270/27

= 10

overhead assigned to Job 466 = Direct labor hours x Predetermined overhead rate

= 10 x 24

= $240

Kindly comment if you need further assistance. Thanks